# Basis Construction for Spline Spaces over Arbitrary Partitions from a Dimensional Stable Perspective

### v1

@@ 3.2 Two-Dimensional Case
    implying $s\in\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}$. Hence, $S_{d}^{\mu}(\mathscr{T})\subseteq\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}$.
    Conversely, under the EEE conditions, $\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\in S_{d}^{\mu}(\mathscr{T})$ for $j=1,2,\ldots,h$, so $\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}\subseteq S_{d}^{\mu}(\mathscr{T})$. Therefore,
    $S_{d}^{\mu}(\mathscr{T})=\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}.$
@@
    and since $\{N_{1},N_{2},\ldots,N_{n}\}$ is a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}(\mathscr{T}))$, it follows that
@@
In conclusion, $S_{d}^{\mu}(\mathscr{T})=\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}$, and the functions in $\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}$ are linearly independent. Thus, these functions form a basis for $S_{d}^{\mu}(\mathscr{T})$, and consequently, $h=\dim S_{d}^{\mu}(\mathscr{T})$.
$\Box$
The construction above, termed semi-explicit, combines basis functions over $\mathrm{ext}_{s}(\Delta)$ with the EEE conditions and is detailed in Section 4.
In the following, based on the concept of Construction-Suitable Partitions and EEE conditions, we provide the framework for the construction of basis functions for the spline space over arbitrary polygonal partitions. Before presenting the framework, we establish a lemma that ensures the existence of a compatible pair $(\mathrm{ext}_{s}(\Delta),\mathcal{M})$.
Lemma 3.7
@@
Proof  Select the cross-cut partition $\mathrm{ext}_{s}(\Delta)$ extended from $\Delta$ and adopt the basis construction method $\mathcal{M}$ for the spline space over cross-cut partitions, as described in Wang (2013). This ensures that, for any polygonal partition $\Delta$, a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$ can always be constructed.
$\Box$
@@
Figure 3 illustrates the basis construction process, where CSP denotes Construction-Suitable Partition. The following Algorithm 1 outlines the framework for constructing a basis for the polynomial spline space $S_{d}^{\mu}(\Delta)$.
@@ 4.1 Comparisons
This subsection compares the proposed framework with other basis construction methods. First, we compare our method with the knot insertion and the Oslo algorithm to highlight its novelty. Second, using spline space over T-mesh as an example, we conduct a comprehensive comparison of basis construction methods, classifying them into three categories: explicit, implicit, and semi-explicit constructions. The proposed framework is semi-explicit. Additionally, we briefly compare spline basis construction with finite element basis construction to clarify their distinctions.
@@ 4.1.1 A Comparison of One-Dimensional Case
Table 2 summarizes the differences between the Oslo algorithm and our method in the one-dimensional case.
@@ 4.1.2 A Comparative Analysis of Basis Construction
We classify spline space basis function constructions into three categories: explicit, semi-explicit, and implicit. Using the spline space over a T-mesh as an example, we compare their advantages and disadvantages. Additionally, we briefly contrast spline basis construction with finite element basis construction to highlight the lower degrees of freedom in spline methods.
Definition 4.1
@@
  • Semi-Explicit Construction: Basis functions of $S_{d}^{\mu}(\Delta)$ are defined explicitly but require solving a linear system for parameter selection.
@@
Based on Definition 4.1, the method proposed in this paper is classified as a semi-explicit construction. It derives basis functions for the spline space by solving a linear system defined by the EEE condition, using combination coefficients from a set of basis functions in the solution space. This subsection examines the advantages and limitations of explicit, semi-explicit, and implicit construction methods through examples.
Example 4.1
@@
  • Semi-Explicit Construction: The proposed method introduces the first semi-explicit construction, deriving basis functions for spline spaces on arbitrary polygonal partitions. It combines an explicit construction with EEE conditions to ensure a complete basis. The explicit component must provide a basis, not merely linearly independent generators.
@@
Table 3 compares explicit, semi-explicit, and implicit construction methods for spline spaces, focusing on construction approach, theoretical robustness, mesh refinement flexibility, computational complexity, and applicability to theoretical and practical contexts.
Aspect Explicit Construction Semi-Explicit Construction Implicit Construction
@@
Explicit construction is simple but often produces linearly dependent functions, failing to form a complete basis and requiring restrictive mesh conditions for refinement. Semi-explicit construction addresses these issues through EEE conditions, supporting flexible refinement on arbitrary meshes, though computing these conditions is computationally intensive, highlighting a key area for future optimization. Implicit construction is highly effective for practical applications, such as isogeometric analysis, due to its straightforward handling of smoothness constraints, but it offers minimal theoretical advancements.
@@ 4.2 Examples
  • Verify that $(\mathrm{ext}_{3}(\Delta_{\mathrm{MS}}),\mathcal{M})$ is compatible, then construct basis functions for $S_{2}^{1}(\mathrm{ext}_{3}(\Delta_{\mathrm{MS}}))$ using the method in (Wang, 2013):
@@
    $\frac{\partial^{2}}{\partial\boldsymbol{n}^{2}}f(x,y)=a^{2}\frac{\partial^{2}f}{\partial x^{2}}+2ab\frac{\partial^{2}f}{\partial x\partial y}+b^{2}\frac{\partial^{2}f}{\partial y^{2}}.$ (4.10)
@@
    $\frac{\partial^{2}}{\partial\boldsymbol{n}^{2}}[L_{i}]^{2}=2(\boldsymbol{n}_{i}\boldsymbol{n}^{\top})^{2}.$ (4.11)
@@
      For $N_{7}=p_{1}[L_{FG}]_{+}^{2}+p_{\text{FC}}[L_{\text{FC}}]_{+}^{2}-p_{\text{FE}}[L_{\text{EI}}]_{+}^{2}$, from Figure 6, $L_{\text{EI}}]_{+}^{2}=[L_{II^{\prime}}]_{+}^{2}$ and $L_{\text{FC}}]_{+}^{2}=[L_{CC^{\prime\prime}}]_{+}^{2}$ are $C^{\infty}$ in $GF$ except at $F$. Since $N_{i}\in C^{1}(\Delta)$, $i=1,\ldots,9$, and $\sum c_{i}N_{i}$ is $C^{1}$ at interior vertices of $\Delta$, edge $FG$ (except $F$) requires $C^{1}$ continuity.
@@
    $\begin{pmatrix}-p_{1}&0&2p_{\text{DF}}\\ p_{\text{FE}}&-p_{2}&0\\ 0&2p_{\text{ED}}&-p_{3}\end{pmatrix}\begin{pmatrix}c_{7}\\ c_{8}\\ c_{9}\end{pmatrix}=\begin{pmatrix}0\\ 0\\ 0\end{pmatrix}$. (4.12)
@@
      If $p_{1}p_{2}p_{3}=4p_{\text{ED}}p_{\text{DF}}p_{\text{FE}}$, then $\mathrm{rank}(M)=2$, allowing a solution $(c_{7},c_{8},c_{9})=(c_{7},0,0)$. The basis for $S_{2}^{1}(\Delta)$ is $\{N_{i}\}_{i=1}^{7}$. Interior edges, such as $DA$, are degenerate. The basis function $N_{7}$ satisfies the EEE conditions, indicating automatic degeneracy of the extended edges $FG$, EIEI, and $HD$.
We construct a basis for the spline space over the Morgan-Scott partition. Despite mild constraints on the partition, the method extends to any Morgan-Scott partition for basis construction. Edge cofactors in the Morgan-Scott partition are typically degenerate. Non-degenerate edge cofactors exist only when specific geometric conditions are met.
We summarize the construction of a basis for the spline space on the Morgan-Scott partition as follows:
  • Basis functions can be constructed over any dimensional unstable partitions, with instability characterized by the EEE conditions.
  • Dimensional instability occurs not only with non-degenerate smooth edge cofactors but also in partitions with some degenerate smooth edge cofactors. The Morgan-Scott partition is a typical example.
@@ 5 Conclusion and future work
This study presents a framework for constructing basis functions for spline spaces over arbitrary meshes. We first address limitations of knot insertion algorithms in one-dimensional cases by analyzing properties of refined and coarse spline spaces, establishing a general framework. This framework is extended to two-dimensional polygonal partitions. We introduce the concept of construction-suitable partitions, as defined in Definition 3.3, and clarify the relationship between dimensional stability and basis function construction: basis functions require a dimensional stable mesh. We define compatibility between a polygonal mesh $\Delta$} and a basis construction method $\mathcal{M}$, showing that an extended mesh $\mathrm{ext}_{s}({\Delta})$ (penetrating or quasi-penetrating) is compatible with the method. Using the EEE condition from Zhong et al. (2025), we develop Algorithm 1, a general framework for two-dimensional basis construction, demonstrating that basis functions can be constructed on dimensionally unstable meshes, challenging prior assumptions.
We classify spline basis construction methods into three categories—explicit, semi-explicit, and implicit—based on their basis function representations, and provide a comparative analysis of their strengths and limitations. We also distinguish spline basis construction from finite element basis construction, noting that splines offer fewer degrees of freedom, suggesting potential as an alternative to traditional finite element methods. Applying our framework to the Morgan-Scott partition, we construct a basis and show that it exhibits degenerate smooth edge cofactors, confirming dimensional instability in both non-degenerate and degenerate meshes, as discussed in Section 4. This provides a valuable case study for further investigation of dimensional instability.
@@
  • Optimizing basis functions: The semi-explicit method yields basis functions with coefficients satisfying the EEE condition. Selecting coefficients to enhance basis function properties is a key question for future exploration.
6 Acknowledgement
We acknowledge Dr. Shicong Zhong, School of Mathematical Sciences, University of Science and Technology of China, under the supervision of Prof. Falai Chen, for providing the EEE conditions Zhong et al. (2025), referred to as Zhong conditions, which are central to this study. We also thank the anonymous reviewers for their valuable feedback and suggestions.

### v2

@@ 3.2 Two-Dimensional Case
    implying $s\in\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}$. Hence, $S_{d}^{\mu}(\Delta)\subseteq\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}$.
    Conversely, under the EEE conditions, $\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\in S_{d}^{\mu}(\Delta)$ for $j=1,2,\ldots,h$, so
    $\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}\subseteq S_{d}^{\mu}(\Delta)$
    Therefore,
    $S_{d}^{\mu}(\Delta)=\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}.$
@@
    and since $\{N_{1},N_{2},\ldots,N_{n}\}$ is a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$, it follows that
@@
In conclusion, $S_{d}^{\mu}(\Delta)=\mathrm{span}\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}$, and the functions in $\left\{\sum_{i=1}^{n}\tilde{c}_{i,j}N_{i}\right\}_{j=1}^{h}$ are linearly independent. Thus, these functions form a basis for $S_{d}^{\mu}(\Delta)$, and consequently, $h=\dim S_{d}^{\mu}(\Delta)$.
$\square$
The construction above, termed semi-implicit, combines basis functions over $\mathrm{ext}_{s}(\Delta)$ with the EEE conditions and is detailed in Section 4.
In the following, based on the concept of Basis Construction-Suitable Partitions and EEE conditions, we provide the framework for the construction of basis functions for the spline space over arbitrary polygonal partitions. Before presenting the framework, we establish a lemma that ensures the existence of a compatible pair $(\mathrm{ext}_{s}(\Delta),\mathcal{M})$.
Lemma 3.21.
@@
Proof 3.22.
Select the cross-cut partition $\mathrm{ext}_{s}(\Delta)$ extended from $\Delta$ and adopt the basis construction method $\mathcal{M}$ for the spline space over cross-cut partitions, as described in Wang (2013). This ensures that, for any polygonal partition $\Delta$, a basis for $S_{d}^{\mu}(\mathrm{ext}_{s}(\Delta))$ can always be constructed.
$\square$
@@
Figure 3 illustrates the basis construction process, where CSP denotes Basis Construction-Suitable Partition. The following Algorithm 1 outlines the framework for constructing a basis for the polynomial spline space $S_{d}^{\mu}(\Delta)$.
@@
Although Definition 3.15 of the EEE condition requires that all points on an edge $l$, including its endpoints, satisfy the condition, Algorithm 1 constructs basis functions for the spline space $S_{d}^{\mu}(\Delta)$, where the $C^{k}$ smoothness condition is inherently satisfied at all interior vertices of $\Delta$. Consequently, the EEE condition can be relaxed to apply only to the edges, excluding their interior vertices.
@@ 4.1 Comparisons
This subsection compares the proposed framework with other basis construction methods. First, we compare our method with the knot insertion and the Oslo algorithm to highlight its novelty. Second, using spline space over T-mesh as an example, we conduct a comprehensive comparison of basis construction methods, classifying them into three categories: explicit, implicit, and semi-implicit constructions. The proposed framework is semi-implicit. Additionally, we briefly compare spline basis construction with finite element basis construction to clarify their distinctions.
@@ 4.1.1 A Comparison of One-Dimensional Case
Table 2 presents the differences between the Oslo algorithm and our method in the one-dimensional case. In fact, by selecting the columns of the transition matrix $A$ as a set of particular solutions to the Extended Edge Elimination (EEE) condition, which inherently satisfy the condition, our method constructs a set of normalized B-splines in $S_{d}^{\mu}(\Delta_{2})$. Consequently, our algorithm can be regarded as a generalization of the Oslo algorithm, representing a specific set of solutions to the EEE condition.
@@ 4.1.2 A Comparative Analysis of Basis Construction
We classify spline space basis function constructions into three categories: explicit, semi-implicit, and implicit. Using the spline space over a T-mesh as an example, we compare their advantages and disadvantages. Additionally, we briefly contrast spline basis construction with finite element basis construction to highlight the lower degrees of freedom in spline methods.
Definition 4.1.
@@
  • Semi-Implicit Construction: Basis functions of $S_{d}^{\mu}(\Delta)$ are defined explicitly but require solving a linear system for parameter selection.
@@
Based on Definition 4.1, the method proposed in this paper is classified as a semi-implicit construction. It derives basis functions for the spline space by solving a linear system defined by the EEE condition, using combination coefficients from a set of basis functions in the solution space. This subsection examines the advantages and limitations of explicit, semi-implicit, and implicit construction methods through examples.
Example 4.2.
@@
  • semi-implicit Construction: The proposed method introduces the first semi-implicit construction, deriving basis functions for spline spaces on arbitrary polygonal partitions. It combines an explicit construction with EEE conditions to ensure a complete basis. The explicit component must provide a basis, not merely linearly independent generators.
@@
Table 3 compares explicit, semi-implicit, and implicit construction methods for spline spaces, focusing on construction approach, theoretical robustness, mesh refinement flexibility, computational complexity, and applicability to theoretical and practical contexts.
Aspect Explicit Construction semi-implicit Construction Implicit Construction
@@
Explicit construction is simple but often produces linearly dependent functions, failing to form a complete basis and requiring restrictive mesh conditions for refinement. semi-implicit construction addresses these issues through EEE conditions, supporting flexible refinement on arbitrary meshes, though computing these conditions is computationally intensive, highlighting a key area for future optimization. Implicit construction is highly effective for practical applications, such as isogeometric analysis, due to its straightforward handling of smoothness constraints, but it offers minimal theoretical advancements.
@@ 4.2 Examples
  • Verify that $(\mathrm{ext}_{3}(\Delta_{\mathrm{MS}}),\mathcal{M})$ is compatible, then construct basis functions for $S_{2}^{1}(\mathrm{ext}_{3}(\Delta_{\mathrm{MS}}))$ using the method in Wang (2013):
@@
    (4.10) $\frac{\partial^{2}}{\partial\boldsymbol{n}^{2}}f(x,y)=a^{2}\frac{\partial^{2}f}{\partial x^{2}}+2ab\frac{\partial^{2}f}{\partial x\partial y}+b^{2}\frac{\partial^{2}f}{\partial y^{2}}.$
@@
    (4.11) $\frac{\partial^{2}}{\partial\boldsymbol{n}^{2}}[L_{i}]^{2}=2(\boldsymbol{n}_{i}\boldsymbol{n}^{\top})^{2}.$
@@
      For $N_{7}=p_{1}[L_{FG}]_{+}^{2}+p_{\text{FC}}[L_{\text{FC}}]_{+}^{2}-p_{\text{FE}}[L_{\text{EI}}]_{+}^{2}$, from Figure 6, $L_{\text{EI}}]_{+}^{2}=[L_{II^{\prime}}]_{+}^{2}$ and $L_{\text{FC}}]_{+}^{2}=[L_{CC^{\prime\prime}}]_{+}^{2}$ are $C^{\infty}$ in $GF$ except at $F$. Since $N_{i}\in C^{1}(\Delta)$, $i=1,\ldots,9$, then $\sum c_{i}N_{i}$ is $C^{1}$ at $F$, then edge $FG$ (except $F$) requires $C^{2}$ continuity.
@@
    (4.12) $\begin{pmatrix}-p_{1}&0&2p_{\text{DF}}\\ p_{\text{FE}}&-p_{2}&0\\ 0&2p_{\text{ED}}&-p_{3}\end{pmatrix}\begin{pmatrix}c_{7}\\ c_{8}\\ c_{9}\end{pmatrix}=\begin{pmatrix}0\\ 0\\ 0\end{pmatrix}$.
@@
      If $p_{1}p_{2}p_{3}=4p_{\text{ED}}p_{\text{DF}}p_{\text{FE}}$, then $\mathrm{rank}(M)=2$, allowing a solution
      $(c_{7},c_{8},c_{9})=(2p_{2}p_{DF},2p_{DF}p_{FE},p_{1}p_{2}).$
      The basis for $S_{2}^{1}(\Delta)$ is
      $\{N_{i}\}_{i=1}^{6}\cup\{2p_{2}p_{DF}\cdot N_{7}+2p_{DF}p_{FE}\cdot N_{8}+p_{1}p_{2}\cdot N_{9}\}.$
      The basis function $2p_{2}p_{DF}\cdot N_{7}+2p_{DF}p_{FE}\cdot N_{8}+p_{1}p_{2}\cdot N_{9}$ satisfies the EEE conditions, indicating automatic degeneracy of the extended edges $FG$, EIEI, and $HD$.
This example constructs basis functions for the spline space over the Morgan-Scott partition, applicable to any such partition despite mild geometric constraints. We show that dimensional instability stems from the rank deficiency of the coefficient matrix in the Extended Edge Elimination (EEE) conditions, which is determined by edge cofactors associated with specific geometric configurations of the partition. This rank deficiency causes basis functions over the extended quasi-cross-cut partition to form linear combinations that produce additional basis functions satisfying the EEE conditions, thereby increasing the total number of basis functions. The basis construction is summarized as follows:
  • Basis functions can be constructed for dimensional unstable partitions, with instability defined by the EEE conditions.
  • Dimensional instability results from degenerate EEE conditions.
  • Instability occurs in partitions with both non-degenerate and degenerate smooth edge cofactors, as exemplified by the Morgan-Scott partition.
@@
Finally, we derive a dimension formula for the spline space $S_{d}^{\mu}(\Delta)$ over an arbitrary partition $\Delta$ of a domain $\Omega$, utilizing the EEE conditions.
Theorem 4.6.
Let $\Delta$ be a partition of $\Omega$, and let $\mathrm{ext}_{s}(\Delta)$ be the extended cross-cut partition of $\Delta$ with $L$ cross-cuts and $V$ interior vertices $A_{1},\dots,A_{V}$ in $\Omega$, where $N_{i}$ cross-cuts intersect at $A_{i}$ for $i=1,\dots,V$. The dimension of the spline space $S_{d}^{\mu}(\Delta)$ is given by
$\dim S_{d}^{\mu}(\Delta)=\binom{d+2}{2}+L\binom{d-\mu-1}{2}+\sum_{i=1}^{V}k_{d}^{\mu}(N_{i})-\mathrm{rank}(M),$
where $M$ is the coefficient matrix of the EEE conditions.
Theorem 4.7.
Let $\Delta$ be a partition of $\Omega$, and let $\mathrm{ext}_{s}(\Delta)$ be the extended quasi-cross-cut partition of $\Delta$ with $L_{2}$ rays and $V$ interior vertices $A_{1},\dots,A_{V}$ in $\Omega$, where $N_{i}$ cross-cuts and rays intersect at $A_{i}$ for $i=1,\dots,V$. The dimension of the spline space $S_{d}^{\mu}(\Delta)$ is given by
$\dim S_{d}^{\mu}(\Delta)=\binom{d+2}{2}+L_{2}\binom{d-\mu-1}{2}+\sum_{i=1}^{V}k_{d}^{\mu}(N_{i})-\mathrm{rank}(M),$
where $M$ is the coefficient matrix of the EEE conditions.
@@ 5 Conclusion and future work
This study presents a framework for constructing basis functions for spline spaces over arbitrary meshes. We first address limitations of knot insertion algorithms in one-dimensional cases by analyzing properties of refined and coarse spline spaces, establishing a general framework. This framework is extended to two-dimensional polygonal partitions. We introduce the concept of Basis Construction-Suitable partitions, as defined in Definition 3.10, and clarify the relationship between dimensional stability and basis function construction: basis functions require a dimensional stable mesh. We define compatibility between a polygonal mesh $\Delta$} and a basis construction method $\mathcal{M}$, showing that an extended mesh $\mathrm{ext}_{s}({\Delta})$ (penetrating or quasi-penetrating) is compatible with the method. Using the EEE condition from Zhong et al. (2025), we develop Algorithm 1, a general framework for two-dimensional basis construction, demonstrating that basis functions can be constructed on dimensionally unstable meshes, challenging prior assumptions.
We classify spline basis construction methods into three categories—explicit, semi-implicit, and implicit—based on their basis function representations, and provide a comparative analysis of their strengths and limitations. We also distinguish spline basis construction from finite element basis construction, noting that splines offer fewer degrees of freedom, suggesting potential as an alternative to traditional finite element methods. Applying our framework to the Morgan-Scott partition, we construct a basis and show that it exhibits degenerate smooth edge cofactors, confirming dimensional instability in both non-degenerate and degenerate meshes, as discussed in Section 4. This provides a valuable case study for further investigation of dimensional instability.
Finally, we would like to emphasize that this work is not only advances the theoretical understanding of spline spaces but also enhances adaptive refinement applications. By targeting only cells not meeting the error threshold $\epsilon$ for refinement, our framework avoids refining cells that satisfy $\epsilon$. This ensures error reduction in targeted cells without affecting those meeting the threshold, minimizing unnecessary computations.
@@
  • Optimizing basis functions: The semi-implicit method yields basis functions with coefficients satisfying the EEE condition. Selecting coefficients to enhance basis function properties is a key question for future exploration.